Write the slope-intercept equation for the line that passes through (10, -4) and has slope -3/10. Please show all of your work.
Answer by mananth(16946)

m=-0.3

Plug value of the slope and point (10 ,-4) in
Y =mx+b
-4.00=-3 +b
b=-4+3
b= -1
So the equation will be
Y =-0.3 x-1